ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

19. History of Mathematics

1991 MS Classification: 1 Barrow-Green, June, The Open University, UK Mathematics in Britain 1860-1940: a database for the World Wide Web An extensive database has been constructed giving very full details of mathematical activity in Britain from 1860 to 1940. It will soon be available on the World Wide Web, thus making accessible a considerable amount of information to support the growing interest in the history of mathematics over this period. In particular, the growth of research (eg the schools of GH Hardy and HF Baker), the emergence of the mathematical profession, changes in the mathematical syllabus at universities, and ways in which these aspects of mathematics were related, can now be investigated more thoroughly than before. In this presentation, the contents of the database, its design, its transference to the Web and its use in supporting research will be described.